OpenStudy (anonymous):

Please show work what is 25 % of 980 It would mean a lot if you would help :) Thank you :3

OpenStudy (anonymous):

Can you help :( 

OpenStudy (anonymous):

980/100 = 9.8 9.8*25 = 245

OpenStudy (anonymous):

or 980 * 0.25 = 245

OpenStudy (anonymous):

What is 25% of 980? What equals 25% of 980 What equals 25 percent of 980 What equals 25 per 100 of 980 \[? = \frac{25}{100} * \frac{980}{1}\]